OpenStudy (livya15):

Find the Derivative (dy/dx) of the following problem:

OpenStudy (livya15):

\[y=\cot ^{2} 5x \sin5x \]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=2\cot 5x \left( -\csc ^25x \right)*5\sin 5x+2\cot ^25x*\cos 5x*5\] \[=10\cot 5x \sin 5x \left( -\csc ^25x+\cot 5x \right)\] ?

OpenStudy (livya15):

What happened to the cot5x and the cot^2 5x? @surjithayer

OpenStudy (anonymous):

Did you get first step?

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=2\cot 5x \left( -cosec ^25x \right) \cdot5\sin 5x+2\cot ^25x \cdot \cos 5x \cdot 5\]

OpenStudy (livya15):

I understand that step, i am just confused of how you simplify it...

OpenStudy (phi):

I would simplify first, then find the derivative cot = cos/sin and cot^2 5x * sin 5x = cos^2 5x / sin^2 5x * sin 5x =cos^2 5x / sin 5x = (1- sin^2 5x) / sin 5x = 1/ sin 5x - sin 5x now take the derivative.

OpenStudy (livya15):

thanks